# Positioning is All You Need

### v1

@@ Abstract
One can drive around safely using a GPS without memorizing a world map (not to mention the dark regions that humans have never explored) because we only pay attention to the next instruction of turning. Such simple observation with attention mechanism has a profound implication on our understanding of how the brain works. The nonlinear effects generated by inhibitory networks have long been known to provide a “winner-take-all” mechanism for localized attention. This attention mechanism, at the system level, implements the selection/positioning operation needed by the hippocampus to provide spatio-temporal context to the neocortex. The localized receptive field of sensory organs, when inspected under the framework of embodied cognition, turns from a constraint to an advantage. Since we only need access to a local map, positioning works better than reconstruction. Such simple intuition implies that the attention mechanism is all you need to understand the flexible behavior generated by hippocampal-neocortical systems. Geometrically, we present a novel manifold positioning framework to explain the principle of localized embodied cognition. Through the co-evolution of the hippocampus and the neocortex, the positioning operation implemented by the attention mechanism can be interpreted as a nonlinear projection linking the discovery of local subspace structure by the neocortex (a sensorimotor machine generating the world map) with the navigation task in mind and $without$ discovering global manifold topology by the hippocampus (the neocortex’s map index).
@@ Introduction
The sensory organs attached to mammalian brains can access only a localized region in space at any specific time. Such spatio-temporal locality of sensory observation has a profound implication for the mental construction of the internal world (a.k.a., internal representation  (?)) including consciousness by the mammalian brain. Nature has discovered a clever solution to learn a global representation of the physical world with a local receptive field of sensory cortex by flexible behavior including spatial navigation. How do mammalian brains encode the spatial information about the physical world into the population of spike sequence in the temporal domain? Such a question is at the holy grail of neuroscience whose importance to cognitive science is equivalent to the relativity theory in modern physics.

### v2

@@ Abstract
One can drive safely with a GPS without memorizing a world map (not to mention the dark regions that humans have never explored). Such a locality-based attention mechanism has a profound implication on our understanding of how the brain works. This paper refines the existing embodied cognition framework by turning the locality from a constraint to an advantage. Analogous to GPS-based navigation, positioning represents a computationally more efficient solution to flexible behaviors than reconstruction. This simple intuition implies that positioning is all you need to understand cognitive functions generated by hippocampal-neocortical systems. That is, the neocortex generates thousands of local maps whose indexing is maintained by the hippocampus. Geometrically, we present a simple manifold positioning framework to explain the principle of localized embodied cognition. The positioning operation implemented by the attention mechanism can be interpreted as a nonlinear projection linking the discovery of local subspace structure by the neocortex (a sensorimotor machine interacting with the world locally) for the navigation task in mind $without$ discovering global manifold topology.
@@ Introduction
Sensory organs attached to the mammalian brain can access only a localized region in space at any specific time. Such spatio-temporal locality of sensory observation has a profound implication for the mental construction of the internal world (a.k.a., internal representation  (?)) including consciousness by the mammalian brain. Nature has discovered a clever solution to learn a global representation of the physical world with a local receptive field of sensory cortex by flexible behavior, including spatial navigation. How do mammalian brains encode the spatial information about the physical world into the population of spike sequences in the temporal domain? Such a question is at the holy grail of neuroscience whose importance to cognitive science is equivalent to relativity theory in modern physics.